two forces are acting at an angle of 120 the bigger force is 40n n and resultant is perpendicular to smaller force, find the smaller force

Grade:12th pass

1 Answers

Arun
4 years ago
Dear Vipul
 
Answer is 20 N
 
Solution-
Tan (α) = psin (θ)/p + qcos (θ).
Alpha = 90, tan90 = 1/0.
θ = 120 = 180 - 120 = 60, sin60 = root3/2.
Cos60 = 1/2.
1/0 = p (root3/2)/p-40 (1/2).
P = 20.